OpenStudy (anonymous):

A phone company charges its customers as follows: For the first 500 minutes used each month, the customer is charged a flat rate of $50. These 500 minutes are "included" in their monthly plan. For any additional minutes used, the customer is charged an additional $.10 per minute. Which type of function would best model this situation: a greatest integer function or a piecewise function that is not a greatest integer function? Explain

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\( f(x) = \begin{cases} 50, & \text{if } x \le 500 \\ 50 + 0.1(x-500), & \text{if }x \gt 500a \end{cases} \)

OpenStudy (anonymous):

@ganeshie8 thank you

ganeshie8 (ganeshie8):

np :) clearly we need a piecewise function that is not a greatest integer function because the phone bill can be in cents also... for example, if u have used 501 minutes, f(x) will be 50+0.1(501-500) = 50.1. which is clearly not an integer. so you need a piecewise function which is not a greatest integer funciton